What would be the interest rate on a 10-year Treasury bond, given the following information? kpr = 2% MR = 0.1% for a 1-year loan, increasing by 0.1% each additional year. LR = 0.5% DR= 0 for a 1-year loan, 0.1% for a 2-year loan, increasing by 0.1% for each additional year. Expected inflation rates: Year 1 = 3.0% Year 2 = 4.0% Year 3 and thereafter: 5.0% 6.7% 9.1% 7.7% 8.9%
Answer #1
r* = 2%
YEAR INFLATION LP MRP DRP AVERAGE INFLATION PER YEAR(IP) K = r*+IP+LP+MRP+DRP
1 3.00% 0.50% 0.10% 0.00% 3.00% 3.00% 5.60%
2 4.00% 0.50% 0.20% 0.10% (3%+4%)/2= 3.50% 6.30%
3 5.00% 0.50% 0.30% 0.20% (3%+4%+5%)/3= 4.00% 7.00%
4 6.70% 0.50% 0.40% 0.30% (3%+4%+5%+6.7%)/4= 4.68% 7.88%
5 9.10% 0.50% 0.50% 0.40% (3%+4%+5%+6.7%+9.1%)/5= 5.56% 8.96%
6 7.70% 0.50% 0.60% 0.50% (3%+4%+5%+6.7%+9.1%+7.7%)/6= 5.92% 9.52%
7 8.90% 0.50% 0.70% 0.60% (3%+4%+5%+6.7%+9.1%+7.7%+8.9)/7= 6.34% 10.14%
8 8.90% 0.50% 0.80% 0.70% (3%+4%+5%+6.7%+9.1%+7.7%+(8.9x2))/8= 6.66% 10.66%
9 8.90% 0.50% 0.90% 0.80% (3%+4%+5%+6.7%+9.1%+7.7%+(8.9x3))/9= 6.91% 11.11%
10 8.90% 0.50% 1.00% 0.90% (3%+4%+5%+6.7%+9.1%+7.7%+(8.9x4))/10= 7.11% 11.51%
Answer is 11.51%